### v1

@@ Abstract
Ring systems have recently been discovered around several small bodies in the outer Solar System through stellar occultations. While such measurements provide key information on ring geometry and dynamical interactions, little is known about their composition, grain-size distribution, origin, lifetime, or evolutionary pathways. Here we report near-infrared observations from the James Webb Space Telescope (JWST) of a stellar occultation by the Centaur (10199) Chariklo, providing unprecedented constraints on the material properties of a small-body ring system and offering insights into their origin and evolution. These measurements reveal that Chariklo’s inner dense ring contains predominantly micrometre-sized particles and exhibits a significant increase in opacity compared to previous observations, suggesting active replenishment events. Most strikingly, the outer ring shows a much weaker near-infrared occultation signature than in earlier visible-light detections. This discrepancy may indicate ongoing material loss, implying that the outer ring is transient, or it may reflect wavelength-dependent opacity consistent with a dusty structure dominated by 0.2–0.5 m silicate grains. These scenarios, not mutually exclusive, point to an unprecedented level of complexity in small-body ring systems, unlike anything observed around other minor bodies in the Solar System.
@@
KBOs, TNOs, Centaurs, stellar occultations, JWST, Chariklo, ringsStellar occultations are a powerful tool in planetary astronomy, offering unique opportunities to probe the properties of distant solar system bodies and their environments with exceptional precision. When a celestial object, such as an asteroid, passes in front of a star, the brief dimming of starlight reveals information about the occulting body’s size, shape, and surroundings [1, 2]. This technique has also enabled the detection and characterisation of tenuous atmospheres [3, 4, 5], or has placed stringent upper limits on their presence [2]. Notably, stellar occultations have led to the discovery and study of ring systems not only around the giant planets [6], but more recently around small bodies, including the Centaurs Chariklo [7] and Chiron [8, 9, 10], the dwarf planet Haumea [11], and the trans-Neptunian object Quaoar [12, 13]. Among these, Chariklo was the first small Solar System body discovered to host a ring system, composed of two sharply confined components, C1R and C2R, orbiting at 390 and 405 km from its centre, respectively [7, 14, 15]. Yet, despite extensive occultation campaigns and multiple studies on ring dynamics [16, 17, 18], the composition, grain-size distribution and origin of the rings have remained poorly constrained and have defied a clear explanation for over a decade. Constraining the composition of rings around small Solar System bodies remains challenging, as their proximity and faintness preclude the isolation of ring spectra using current remote spectroscopy [19]. However, wavelength-dependent optical properties of the grains, which lead to different occultation opacities in the visible and near-infrared, may provide direct constraints on grain properties [20], as it has been proposed for Haumea’s ring system [21].
@@
To investigate the fine structure and composition of Chariklo’s rings beyond the limitations of ground-based observations, we targeted a stellar occultation using the James Webb Space Telescope (JWST). Predicting such an event from a space-based observatory is inherently challenging (see Section 1.1 for details); nevertheless, we successfully predicted and observed a stellar occultation by Chariklo on 18 October 2022. Remarkably, the JWST chord passed just 7.4 km from the body’s centre (Extended Data Figures), allowing us to detect the ring system (see also Section 1.4). No additional dips in brightness were observed that would indicate the presence of other material or orbiting satellites. This observation was conducted simultaneously in two near-infrared bands, centred at 1.5 and 3.2 m (F150W2 and F322W2), including the first occultation detection ever beyond 3 m, which is impossible from the ground. Exactly the same JWST setup was used later in the observations of the occultation event by the ring system of Quaoar [22]. Although the radial locations of both Chariklo’s rings in 2022 remain consistent with earlier measurements, they also exhibit significant variations compared to previous observations. For clarity, we adopt the terms first contact ( contact) and second contact ( contact) to refer to the two ring-plane crossings, each comprising an ingress and an egress: Ingress1–Egress1 and Ingress2–Egress2, respectively. This terminology will be used throughout the manuscript.
@@
Chariklo’s inner dense ring, C1R, is clearly detected in the JWST light curves, exhibiting pronounced diffraction spikes at ingress and egress (see Figure 1), indicative of sharp edges and strong confinement, morphological features reminiscent of Uranus’s rings [23, 24]. Interestingly, C1R is found to be more opaque than before: the pre-JWST measurements provide an average normalized opacity of for this ring, while considering only the JWST measurements gives , indicating a highly significant difference, corresponding to a z-score of z = 4.2 and probability of . In contrast, the outer ring, C2R, displays an unexpectedly weak signal in the JWST measurements. After applying an occultation model with a monolayer abrupt edge, including the effects of Fresnel diffraction (see Section 1.4), C2R was marginally detected at 1 in the F150W2 data, but not in the F322W2 filter, despite both measurements being performed simultaneously and sampling the same ring segment. Compared to previous ground-based visible range observations, this suggests particularly low opacities for C2R.
@@
In the case of a fragmented ring, alternating low- and high-density segments produce azimuthal variability in the observed ring structure, affecting both opacity and equivalent width. While previous occultation measurements yield roughly consistent values for these parameters, the JWST observations reveal significantly different opacity and equivalent width values. Using a simple azimuthal ring density model (see Section 1.5), we estimate that the probability of sampling only low-density (C2R) or high-density (C1R) segments during both ingress and egress in the JWST observations is very low, with p0.004 for C1R and p0.002 for C2R.
@@
Changes in the effective cross-section in occultation measurement can be best described by the equivalent width (), which combines opacity, radial width, and the sensitivity and time resolution of the instrument. is proportional to the amount of material responsible for the extinction within the occulted segment (see Section 1.4 for its formal definition). We carried out a dedicated study of the C2R ring, combining the 2022 JWST occultation measurements (e.g., opacities, equivalent widths) with all available ground-based data from 2013, 2014, and 2017 (see Extended Data Tables and Section 1.4 for details). While no clear trend is seen in between 2013 and 2017, a marked decrease is observed between 2017 and 2022 for C2R (Figure 2, Panel (b)), potentially revealing a decade-long temporal evolution. This apparent fading of the ring may indicate that C2R is actively dispersing. Such depletion could result from collisions among particles, producing smaller grains that are more easily lost. Without effective confinement, these grains are subject to increasing orbital eccentricities due to radiation pressure, leading their orbits to intersect the inner ring [25]. Several known dusty rings in the Solar System exhibit temporal variability in opacity, morphology, or confinement, such as Saturn’s D, E, and F rings, Neptune’s Adams arcs, and Uranus’s ring. For instance, Saturn’s D ring has shown measurable radial shrinkage and brightness changes on decadal timescales, likely driven by micrometeoroid bombardment or electromagnetic forces [26]. The E ring, sourced by Enceladus’s plumes, varies with the moon’s activity and solar cycle [27]. Neptune’s Adams arcs display pronounced morphological evolution and changing brightness over months to years, pointing to a dynamically unstable configuration [28]. Uranus’s ring also shows opacity changes possibly linked to variable confinement mechanisms [29]. These examples illustrate that dusty, low-mass rings can undergo rapid and significant evolution, reinforcing the interpretation of Chariklo’s outer ring as a transient, evolving structure. Strikingly, the JWST opacity and equivalent width of C1R shows an opposite trend, a considerable increase (see Panel (a) in Figure 2, Extended Data Tables, and also [14, 15]). However, the change of in C1R () is approximately 10 times greater than that in C2R (), leaving further questions on the origin of the increase. In C1R, this could be due either to an overall increase in the amount of material or to the rise in the effective cross-section of grains, for example, through collisions between larger particles or rocks inside the ring, producing a cascade of smaller grains.
Since JWST probes a wavelength range where stellar occultations have rarely been observed, the discrepancy with visible-light data may also reflect the optical properties of the ring particles—specifically, a dominance of sub-micron grains whose scattering efficiency decreases at longer wavelengths. Even in the pre-JWST observations, a possible sign of wavelength dependence is the significantly different opacities of the 2013 and 2017 occultation measurements of C2R, using the Danish Z (1 m) and Danish Visual (0.55 m) bands (see Figure 3). To assess whether these variations reflect a genuine wavelength dependence, we adopt a radiative transfer approach using various grain types (e.g., water ice, olivine, carbon) and particle sizes, and compare the resulting extinction values with all available C2R occultation data (see Extended Data Tables and Extended Data Tables) to constrain plausible compositions consistent with the observations. A detailed description of this modelling can be found in Section 1.7. We find that for C2R, a wavelength dependence driven by different optical properties of the ring particles cannot be ruled out. Models dominated by silicates over water reproduce the detection in the F150W2 filter and the non-detection in the F322W2 filter, with pyroxene-rich compositions fitting both the visible and near-infrared opacities when grain sizes lie between 0.2 and 0.5 m (Figure 3). However, we emphasise that these compositional inferences depend on the assumed grain-size distribution and should not be viewed as uniquely limiting the ring’s material composition. In any case, if C2R is indeed dominated by micron-sized dust, its population may be analogous to that of Uranus’s ring, which consists predominantly of similar grains [30]. The ring likely originates from collisions involving a moonlet or a belt of larger parent bodies, potentially metres to kilometres in size, that generate fine particulate debris [31]. Similar dusty ring structures are seen co-orbiting with small Saturnian moons such as Janus, Anthe, Methone, and Pallene, where dust is believed to originate either from the satellites themselves or from co-orbiting larger bodies [32]. A comparable dusty ring structure has also been proposed around Makemake [25].
@@
To explore the wavelength-dependent opacity scenario for C1R, we used a similar radiative transfer calculation, composed of a mixture of two materials with different compositions and potentially different grain sizes (see Section 1.7 and Figure 5). The pre-JWST data, including the 2017 VLT/Ks measurement at 2.1 m, can be well-fitted with mixtures dominated by micrometre-sized water ice or silicate grains with the contribution of very small grains, also indicating a nearly constant opacity throughout the near-infrared. However, when the JWST data are included, it is not possible to find a solution (i.e., a combination of two materials and grain sizes) that simultaneously fits all datasets, particularly the JWST F150W2, F322W2, and the VLT/Ks bands. The best-fit curves in this case contain predominantly silicates with 1–10 m grain sizes, but the overall fits are poor. This clearly indicates that the observed change in opacity cannot be solely attributed to wavelength dependence, and leaves us with the most likely scenario that the opacities in C1R have indeed increased.
@@
The small particle sizes inferred for C2R raise questions about its long-term stability. In general, radiation pressure could expel micron-sized grains on relatively short timescales, potentially leading to structural distortion or dispersal [25]. Yet, occultation data spanning nearly a decade reveal no significant changes in the ring’s radial position (see Table 1), suggesting that a stabilising or replenishing mechanism must be active. Such mechanisms may mirror those proposed for other dusty rings and could involve interactions with C1R. The confinement of C1R is thought to result from the strong 1:3 spin-orbit resonance with Chariklo’s rotation, in which a ring particle completes one orbit for every three rotations of the central body [33]. While effective for C1R, this resonance drives an outward migration of the ring, ultimately risking its departure from the resonance and subsequent dispersion. A small external shepherd moon has been hypothesised to prevent such drift and maintain C1R’s location [33]. If present, this moonlet might share its orbit with C2R and act as a source of dust, analogous to the small satellites observed in Saturn’s and Uranus’s ring systems. This scenario is further supported by our estimate of C2R’s dynamical spreading timescale: based on JWST measurements, we find a median value of only 0.36 years (1 range: 0.23–0.77 years; see Methods, Section 1.6), implying that without confinement or replenishment, the ring should have dispersed since its discovery in 2013. A co-orbiting shepherd satellite could thus play a dual role—maintaining the ring’s sharp edges while continuously supplying fresh material [17]. However, another recent dynamical study about the effect of solar radiation pressure on the stability of a small-body rings also shows that for highly tilted rings, such as around Chariklo, there are two plausible particle size range that can be stable: m and m, continuing to allow for both possible options [18].
The emerging picture suggests that the two scenarios for C2R—rapid erosion and a dusty, micron-scale particle population—are not mutually exclusive. A ring composed of sub-micron grains may be inherently short-lived, as such particles are particularly susceptible to dynamical processes that promote dispersal. We therefore consider both effects jointly, as they may together account for the observed decline in opacity and signal strength of the C2R ring. If the outer ring is indeed degrading over a timescale of just a few years, this would mark it as a transient structure, potentially representative of a broader class of ephemeral rings among small bodies, as hinted by episodic activity and temporary bound structures around another Centaur, Chiron [10]. If instead the C2R ring is linked to Chariklo’s activity as a Centaur-like object, we would again expect small grains, as cometary dust is predominantly sub-micron in size [34, 35]. Crucially, a single occultation in visible light of a sufficiently bright star could distinguish between these scenarios: if the opacity matches values from 2013–2017, grain-size effects are favoured; but if the occultation yields significantly lower opacities or no detection at all, it would instead support the interpretation of an actively fading ring. Either outcome would raise fundamental questions about the longevity of dusty rings and the fate of metre-sized material subject to slow but persistent erosion over decadal timescales. Ultimately, our JWST observations unveil a new and unexpected level of complexity in minor body ring systems—one that challenges existing paradigms and opens an unprecedented window into the transient architectures sculpting the outer Solar System. These findings provide the first evidence of such episodic changes in these seemingly frequent systems, revealing a dynamical behaviour that had remained entirely hidden until now.
@@
Note: : Transmittance; Mid-time: midpoint time of each occultation contact; : radial distance in the ring plane (km); : radial width in the ring plane (km); : apparent opacity; : normal opacity; : apparent optical depth; : normal optical depth; : equivalent width (km); : equivalent depth (km). C2R was not detected in the F322W2 filter due to a low signal-to-noise ratio. 1st and 2nd contact refer to before and after the closest approach, respectively. Values are presented with uncertainties: asymmetric uncertainties are indicated with upper and lower limits (), while symmetric uncertainties are shown with a single value following the symbol.

### v2

@@ Abstract
Ring systems have been discovered around several small bodies in the outer Solar System through stellar occultations. While such measurements provide key information about ring geometry and dynamical interactions, little is known about their origins, lifetimes, evolutionary pathways, or compositions. Here we report near-infrared observations with the James Webb Space Telescope (JWST) of a stellar occultation by (10199) Chariklo, a Centaur known to host a double-ring system. Our JWST measurements show that Chariklo’s inner dense ring has become significantly more opaque than in previous observations, pointing to ongoing replenishment processes or dynamical restructuring. In contrast, the outer ring exhibits a much weaker near-infrared occultation signature than seen in earlier visible-light detections. This discrepancy may reflect material loss, suggesting that the outer ring could be transient, or may arise from wavelength-dependent opacity. These scenarios, which are not mutually exclusive, point to an unprecedented level of complexity in small-body ring systems, distinct from those observed around any other minor bodies in the Solar System.
@@
KBOs, TNOs, Centaurs, stellar occultations, JWST, Chariklo, ringsStellar occultations are a powerful tool in planetary astronomy, offering unique opportunities to probe the properties of distant solar system bodies and their environments with exceptional precision. When a celestial object, such as an asteroid, passes in front of a star, the brief dimming of starlight reveals information about the occulting body’s size, shape, and surroundings [Sicardy2011, Sicardy2006]. This technique has also enabled the detection and characterisation of tenuous atmospheres [Sicardy2016, Sicardy2021, MarquesOliveira2022], or has placed stringent upper limits on their presence [Sicardy2006]. Notably, stellar occultations have led to the discovery and study of ring systems not only around the giant planets [Sicardy2024], but more recently around small bodies, including the Centaurs Chariklo [Braga-Ribas2014] and Chiron [Ortiz2015, Sickafoose2020, Ortiz2023], the dwarf planet Haumea [Ortiz2017], and the trans-Neptunian object Quaoar [Morgado2023, Pereira2023]. Among these, Chariklo was the first small Solar System body discovered to host a ring system, composed of two sharply confined components, C1R and C2R, orbiting at 390 and 405 km from its centre, respectively [Braga-Ribas2014, Berard2017, Morgado2021]. Yet, despite extensive occultation campaigns and multiple studies on ring dynamics [Sicardy2019, Sickafoose2024, Regaly2025], the origin, composition and grain-size distribution of the rings have remained poorly constrained and have defied a clear explanation for over a decade. Although wavelength-dependent optical properties of the grains, which lead to different occultation opacities in the visible and near-infrared, may provide direct constraints on grain properties [bohren2008absorption, Kalup2024], due to the limited number of well-resolved observations and the possible structural changes in the ring, it still remains challenging.
@@
To investigate the fine structure and composition of Chariklo’s rings beyond the limitations of ground-based observations, we targeted a stellar occultation using the James Webb Space Telescope (JWST). Predicting such an event from a space-based observatory is inherently challenging (see Section 1.1 for details); nevertheless, we successfully predicted and observed a stellar occultation by Chariklo on 18 October 2022. Remarkably, the JWST chord passed just 7.4 km from the body’s centre (Extended Data Figures), allowing us to detect the ring system (see also Section 1.4). No additional dips in brightness were observed that would indicate the presence of other material or orbiting satellites. This observation was conducted simultaneously in two near-infrared bands, centred at 1.5 and 3.2 m (F150W2 and F322W2), including the first occultation detection ever beyond 3 m, which is impossible from the ground. Exactly the same JWST setup was used later in the observations of the occultation event by the ring system of Quaoar [Proudfoot2025]. Although the radial locations of both Chariklo’s rings in 2022 remain consistent with earlier measurements, they also exhibit significant variations compared to previous observations.
For clarity, we adopt the terms first contact ( contact) and second contact ( contact) to refer to the two points where the star’s line of sight intersects the ring plane, each comprising an ingress and an egress: Ingress1–Egress1 and Ingress2–Egress2, respectively. This terminology will be used throughout the paper.
Chariklo’s inner dense ring, C1R, is clearly detected in the JWST light curves, exhibiting pronounced diffraction spikes at ingress and egress (see Figure 1), indicative of sharp edges and strong confinement, morphological features reminiscent of Uranus’s rings [Elliot1977, Elliot1987].
@@
Interestingly, C1R is found to be more opaque than before—with all uncertainties throughout this paper reported at the 1- level—the pre-JWST measurements provide an average normal opacity of for this ring, while considering only the JWST measurements gives , indicating a highly significant difference, corresponding to a z-score of z = 4.2 and probability of . In contrast, the outer ring (C2R) shows an unexpectedly weak signal in JWST measurements. After applying an occultation model with a monolayer abrupt edge, including the effects of Fresnel diffraction (see Section 1.4), C2R was marginally detected at 1 in the F150W2 data, but not in the F322W2 filter, despite both measurements being performed simultaneously and sampling the same ring segment. Compared with previous ground-based visible-range observations, this suggests particularly low opacities for C2R.
@@
In the case of a fragmented ring, alternating low- and high-density segments produce azimuthal variability in the observed ring structure. This affects both opacity and equivalent width (), the latter of which combines opacity, radial width, and the instrument’s sensitivity and time resolution (see Section 1.4 for more information). While previous occultation measurements yield roughly consistent values for these parameters, the JWST observations reveal significantly different opacities and equivalent widths. Using a simple azimuthal ring density model, we estimate that the probability of randomly sampling a pair of (both ingress and egress) high-density (C1R) or low-density (C2R) values in the JWST observations against eight other events is very low, with p0.004 for C1R and p0.002 for C2R. See a comprehensive description of this model and calculation in Section 1.5.
@@
Changes in the effective cross-section in occultation measurement are best described by the just-cited equivalent width (). is proportional to the integrated effective cross-section of the grains, responsible for the extinction within the occulted segment (see Section 1.4 for its formal definition). Combining the 2022 JWST occultation measurements with available ground-based data from 2013, 2014, and 2017 (see Extended Data Tables, Extended Data Tables and Section 1.4 for details), we carried out a detailed study of the temporal evolution of the rings. In the case of C2R, while no clear trend is seen in between 2013 and 2017, a marked decrease () is observed between 2017 and 2022 (Figure 2, Panel (b)), potentially revealing a fast evolution in a timescale of a few years. Strikingly, the JWST opacity and equivalent width of C1R shows an opposite trend, a considerable increase (50%, see Panel (a) in Figure 2, Extended Data Tables, and also [Berard2017, Morgado2021]). This change in in C1R () is approximately 10 times greater than that in C2R (), indicating that the material missing from C2R cannot account for the increase in the normal opacity and in C1R. This increase could either be due to an overall increase in the amount of material or to the rise in the effective cross-section of grains, for example, through collisions between larger particles or rocks inside the ring, producing a cascade of smaller grains. The apparent fading of C2R may indicate that it is actively dispersing. Such depletion could result from collisions among particles, producing smaller grains that are more easily lost. Without effective confinement, these grains are subject to increasing orbital eccentricities due to radiation pressure, leading their orbits to intersect the inner ring [Csaba2024]. Several known dusty rings in the Solar System exhibit temporal variability in opacity, morphology, or confinement, such as Saturn’s D, E, and F rings, Neptune’s Adams arcs, and Uranus’s ring. For instance, Saturn’s D ring has shown measurable radial shrinkage and brightness changes on decadal timescales, likely driven by micrometeoroid bombardment or electromagnetic forces [HEDMAN200789]. The E ring, sourced by Enceladus’s plumes, varies with the moon’s activity and the surrounding magnetospheric environment. [KEMPF2008420]. Neptune’s Adams arcs exhibit pronounced morphological evolution and brightness changes over months to years, suggesting a dynamically evolving configuration [renner2014]. Uranus’s ring also shows opacity changes possibly linked to variable confinement mechanisms [FRENCH2012181]. These examples illustrate that dusty, low-mass rings can undergo rapid and significant evolution, reinforcing the interpretation of Chariklo’s outer ring as a transient, evolving structure.
@@
Since JWST probes a wavelength range where stellar occultations have rarely been observed, the discrepancy with visible-light data may also reflect the optical properties of the ring particles—specifically, a dominance of sub-micron grains whose scattering efficiency decreases at longer wavelengths. To assess whether these variations reflect a genuine wavelength dependence, we adopt a radiative transfer approach using various grain types (e.g., water ice, olivine, carbon) and particle sizes, and compare the resulting extinction values with available well-resolved occultation measurements (see Extended Data Tables) to provide a compatibility analysis of any plausible compositions that may be consistent with the observations. A detailed description of this modelling can be found in Section 1.7. For C1R, we used a mixture of two materials with different compositions and potentially different grain sizes (see Section 1.7 and Figure 3). The pre-JWST data, including the 2017 VLT/Ks measurement at 2.1 m, can be well-fitted with mixtures dominated by micrometre-sized water ice or silicate grains with the contribution of tiny grains, also indicating a nearly constant opacity throughout the near-infrared. However, when JWST data are included, no solution (i.e., a combination of two materials and grain sizes) can simultaneously fit all datasets, particularly the JWST F150W2, F322W2, and VLT/Ks bands. The best-fit curves in this case consist predominantly of silicates with grain sizes of 1–10 m, but the overall fits are poor. This clearly indicates that the observed change in opacity cannot be solely attributed to wavelength dependence and leaves us with the most likely scenario: the opacities in C1R have indeed increased. For C2R, it is much harder to perform such an analysis due to the lack of well-resolved occultation profiles. However, as we introduced above, if C2R is actively depleting, this is a strong indication that small grains dominate this ring, as sub-micron grains may be inherently short-lived and particularly susceptible to dynamical processes that promote dispersal.
@@
However, small particle sizes raise questions about long-term stability. In general, radiation pressure could expel micron-sized grains on relatively short timescales, potentially leading to structural distortion or dispersal [Csaba2024]. Yet, occultation data spanning nearly a decade reveal no significant changes in the radial position of both C1R and C2R (see Table 1), suggesting that a stabilising or replenishing mechanism must be active. The confinement of C1R is thought to result from the strong 1:3 spin-orbit resonance with Chariklo’s rotation, in which a ring particle completes one orbit for every three rotations of the central body [Salo2024]. While effective for C1R, this resonance drives an outward migration of the ring, ultimately risking its departure from the resonance and subsequent dispersion. A small external shepherd moon has been hypothesised to prevent such drift and maintain C1R’s location [Salo2024]. If present, this moonlet might share its orbit with C2R and act as a source of ring material, analogous to the small satellites observed in Saturn’s and Uranus’s ring systems. This scenario is further supported by our estimate of C2R’s dynamical spreading timescale: based on JWST measurements, we find a median value of only 0.34 years (1 range: 0.21–0.96 years; see Methods, Section 1.6), implying that without confinement or replenishment, the ring should have dispersed since its discovery in 2013. A co-orbiting shepherd satellite could thus play a dual role—maintaining the ring’s sharp edges while continuously supplying fresh material [Sickafoose2024]. However, another recent dynamical study about the effect of solar radiation pressure on the stability of a small-body rings shows that for highly tilted rings, such as the one around Chariklo, there are two plausible particle size ranges that can be stable: m and m [Regaly2025]. If the outer ring is indeed degrading over a timescale of just a few years, this would mark it as a transient structure, potentially representative of a broader class of ephemeral rings among small bodies, as hinted by episodic activity and temporarily bound structures around another Centaur, Chiron [Ortiz2023].
A single near-future occultation measurement in visible light (different from JWST wavelengths) could confirm the fading of C2R and the strengthening of C1R. Ultimately, our JWST observations unveil a new and unexpected level of complexity in minor body ring systems—one that challenges existing paradigms and opens an unprecedented window into the transient architectures sculpting the outer Solar System. These findings provide the first evidence of such episodic changes in these seemingly frequent systems, revealing a dynamical behaviour that had remained entirely hidden until now.
@@
Note: : Transmittance; Mid-time: midpoint time of each occultation contact; : radial distance in the ring plane (km); : radial width in the ring plane (km); : apparent opacity; : normal opacity; : apparent optical depth; : normal optical depth; : equivalent width (km); : equivalent depth (km). C2R was not detected in the F322W2 filter due to a low signal-to-noise ratio. 1st and 2nd contacts refer to the ones before and after the closest approach, respectively. Values are presented with uncertainties: asymmetric uncertainties are indicated with upper and lower limits (), while symmetric uncertainties are shown with a single value following the symbol.